OpenStudy (anonymous):

A newly discovered planet has a mass 1.5 times compared to that of the earth, and its diameter is 3 times compared to earth. What is the gravitational acceleration on this planet?

OpenStudy (welshfella):

now this is a formula i have definitely forgotten

OpenStudy (anonymous):

Fg = Gm1m2/r^2

OpenStudy (anonymous):

or Fg = Gm/r^2

OpenStudy (welshfella):

oh ok so n1 amd m2 are the masses and r^2 = radius?

OpenStudy (anonymous):

yes ;)

OpenStudy (welshfella):

G is the acceleration due to gravity?

OpenStudy (anonymous):

G = 6.67 X10^-11

OpenStudy (anonymous):

yes :)

OpenStudy (welshfella):

I'm not entirely clear about it and i haven't the time to look it up . sorry

OpenStudy (anonymous):

oh it's okay thank you :)

OpenStudy (anonymous):

can you help me @ganeshie8 :)

OpenStudy (anonymous):

\[mg=G \frac{Mm}{r^2}\] \[g=G \frac{M}{r^2}\]\[g \alpha \frac{M}{r^2}\] \[\frac{g'}{g}=\frac{\frac{M'}{r'^{2}}}{\frac{M}{r^2}}\]\[g'=g \frac{M'r^{2}}{Mr'^{2}}\] M'=Mars of new planet r'=radius of new planet M=mass of earth r=radius of earth Also given \[M'=1.5M\] for the radius also u can calculate